# Consensus over Clustered Networks Using Intermittent and Asynchronous Output Feedback

### v2

@@ II-B Graphs
Let $\mathcal{G}\triangleq(\mathcal{V},\mathcal{E})$ be a graph on $N\in\mathbb{Z}_{\geq 2}$ nodes, where $\mathcal{V}\triangleq[N]$ denotes the node set and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denotes the edge set. If $(p,q)\in\mathcal{E}$ implies $(q,p)\in\mathcal{E}$ for all distinct nodes $p,q\in\mathcal{V}$, then the graph $\mathcal{G}$ is said to be undirected. A path exists between nodes $p,q\in\mathcal{V}$ if there is a sequence of distinct nodes such that $(v_{0}=p,...,v_{k}=q)$ for $k\in\mathbb{Z}_{\geq 0}$, $(v_{s-1},v_{s})\in\mathcal{E}$, and $s\in[k]$. The graph $\mathcal{G}$ is said to be connected if there is a path joining any two distinct nodes in $\mathcal{V}$. The neighbor set of node $p$ is denoted by $\mathcal{N}_{p}\triangleq\{q\in\mathcal{V}\setminus\{p\}:(p,q)\in\mathcal{E}\}$. The adjacency matrix of $\mathcal{G}$ is denoted by $\mathsf{A}\triangleq[a_{pq}]\in\mathbb{R}^{N\times N}$, where $a_{pq}=1$ if and only if $(p,q)\in\mathcal{E}$, and $a_{pq}=0$ otherwise. Self-edges are not employed in this work, that is, $a_{pp}\triangleq 0$ for all $p\in\mathcal{V}$. The degree matrix of $\mathcal{G}$ is denoted by $\Delta\triangleq\text{diag}(\mathsf{A}\cdot 1_{N})\in\mathbb{R}^{N\times N}$. The Laplacian matrix of $\mathcal{G}$ is denoted by $\mathsf{L}\triangleq\Delta-\mathsf{A}\in\mathbb{R}^{N\times N}$. The following result enables the stability analysis provided in Section V.
@@
If $\mathcal{G}$ is static, undirected, and connected, then there exists an orthonormal basis $\beta\triangleq\{v_{1},v_{2},...,v_{N}\}\subset\mathbb{R}^{N}$ for $\text{Range}(\mathsf{L})$ such that $v_{1}=(\sqrt{N}/N)1_{N}$. Consider the matrix $\mathsf{V}\triangleq[v_{2},v_{3},...,v_{N}]\in\mathbb{R}^{N\times N-1}$ and projection $\mathsf{S}\triangleq I_{N}-1_{N}1_{N}^{\top}/N\in\mathbb{R}^{N\times N}$. Then,
@@
The matrix $\mathsf{S}$ is a projection whose image is the orthogonal complement of the agreement subspace $\text{Span}(1_{N})$. That is, for any vector $z\in\mathbb{R}^{N}$, $z$ may be decomposed into the sum of two orthogonal vectors, i.e., $z=z^{\|}+z^{\perp}$, where $z^{\|}\triangleq(1_{N}1_{N}^{\top}/N)z$ and $z^{\perp}\triangleq\mathsf{S}z$. In addition, for any symmetric Laplacian matrix $\mathcal{L}$, one can show by direct computation that $\mathsf{S}\mathcal{L}=\mathcal{L}=\mathcal{L}\mathsf{S}$.
@@
II-C Cluster and Inter-Cluster Sub-Graphs
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a static, undirected, and connected graph. Let $\mathcal{C}(\mathcal{V})\triangleq\{\mathcal{V}_{1},\mathcal{V}_{2},...,\mathcal{V}_{M}\}$ be a fixed partition of the set $\mathcal{V}$, where $M\in[N]$ is some constant. Consequently, $\mathcal{V}_{p}\subseteq\mathcal{V}$ for each $p\in[M]$, $\mathcal{V}_{p}\cap\mathcal{V}_{q}=\varnothing$ for each distinct pair $p,q\in[M]$, and $\cup_{p\in[M]}\mathcal{V}_{p}=\mathcal{V}$. The sets $\mathcal{C}(\mathcal{V})$ and $\mathcal{V}_{q}$, for any $q\in\mathcal{V}$, are referred to as the cluster set of $\mathcal{G}$ and cluster $q$, respectively. Since $\mathcal{C}(\mathcal{V})$ is an arbitrary partition of $\mathcal{V}$, for every $p\in[M]$, cluster $p$ contains at least one node, i.e., $|\mathcal{V}_{p}|\geq 1$. Each cluster induces a sub-graph of $\mathcal{G}$, i.e., for each $p\in[M]$, the induced sub-graph of cluster $p$ is given by $\mathcal{G}[\mathcal{V}_{p}]\triangleq(\mathcal{V}_{p},\mathcal{E}[\mathcal{V}_{p}])$, such that $\mathcal{E}[\mathcal{V}_{p}]\triangleq\{(r,s)\in\mathcal{E}\colon r,s\in\mathcal{V}_{p}\}$. Since $\mathcal{G}$ is undirected, $\mathcal{G}[\mathcal{V}_{p}]$ is undirected. In this work, we only consider cluster sets $\mathcal{C}(\mathcal{V})$ where, for each $p\in[M]$, $\mathcal{G}[\mathcal{V}_{p}]$ is connected. Moreover, since $\mathcal{C}(\mathcal{V})$ is a fixed partition of $\mathcal{V}$, for each $p\in\mathcal{V}$, node $p$ is assigned to one cluster for all time. In particular, for each $p\in\mathcal{V}$, let $\mathcal{C}_{p}$ denote the cluster to which node $p$ belongs, i.e., $p\in\mathcal{C}_{p}$ and $\mathcal{C}_{p}=\mathcal{V}_{q}$ for some $\mathcal{V}_{q}\in\mathcal{C}(\mathcal{V})$. The set of neighbors of node $p$ within $\mathcal{C}_{p}$ is, therefore, given by $\mathcal{N}_{p}\cap\mathcal{C}_{p}$.
@@
For any distinct pair $p,q\in[M]$, let $\mathcal{V}_{pq}\triangleq\mathcal{V}(p,q)\cup\mathcal{V}(q,p)$, such that $\mathcal{V}(p,q)\triangleq\{r\in\mathcal{V}_{p}\colon\exists_{s\in\mathcal{V}_{q}}\ (r,s)\in\mathcal{E}\}$. If $\mathcal{V}_{pq}\neq\varnothing$, then the set $\mathcal{V}_{pq}$ is referred to as inter-cluster $pq$. To simplify the development, we only consider the case where there is at most one inter-cluster between each pair of distinct clusters. The inter-cluster set of $\mathcal{G}$, as determined by $\mathcal{C}(\mathcal{V})$, is denoted by $\mathcal{I}(\mathcal{V})\triangleq\{\mathcal{V}_{pq}\subset\mathcal{V}\colon\forall_{p\neq q\in\mathcal{V}}\ \mathcal{V}_{pq}\neq\varnothing\}$. Since $|\mathcal{C}(\mathcal{V})|=M$ and there can be at most one inter-cluster between each distinct pair of clusters, $|\mathcal{I}(\mathcal{V})|=M^{\ast}\in\{0,1,...,M(M-1)/2\}$. For each distinct pair of $p,q\in[M]$, the sub-graph of $\mathcal{G}$ induced by inter-cluster $pq$ is denoted by $\mathcal{G}[\mathcal{V}_{pq}]\triangleq(\mathcal{V}_{pq},\mathcal{E}_{pq})$, where $\mathcal{E}_{pq}\triangleq\mathcal{E}[\mathcal{V}_{pq}]\setminus(\mathcal{E}[\mathcal{V}_{p}]\cup\mathcal{E}[\mathcal{V}_{q}])$. Further, $\mathcal{G}[\mathcal{V}_{pq}]$ is undirected by construction. Figure 1 illustrates an example of a graph $\mathcal{G}$ that is divided into cluster and inter-cluster sub-graphs. Since $\mathcal{C}(\mathcal{V})$ and $\mathcal{G}$ are fixed, one has that $\mathcal{I}(\mathcal{V})$ is fixed as well. Thus, if node $p\in\mathcal{V}$ belongs to an inter-cluster, then it does so for all time. Moreover, it is possible for a node to belong to multiple inter-clusters simultaneously. The set of neighbors of node $p$ in inter-cluster $rs$ is, therefore, given by $\mathcal{N}_{p}\cap\mathcal{V}_{rs}$.
@@
To facilitate the analysis, $N\times N$ augmented adjacency and Laplacian matrices for a general sub-graph of $\mathcal{G}$ are introduced next. For any non-empty set $S\subseteq\mathcal{V}$, let $\mathsf{A}[S]\triangleq[\mathsf{a}_{pq}]\in\mathbb{R}^{N\times N}$ with
$\mathsf{a}_{pq}\triangleq\Bigg{\{}\begin{aligned} &1,&(p,q)\in\mathcal{E}[S]\\ &0,&\text{otherwise}\end{aligned}$
denote the augmented adjacency matrix of $\mathcal{G}[S]$. Therefore, the augmented adjacency matrix of cluster $p$ and inter-cluster $pq$ are $\mathsf{A}[\mathcal{V}_{p}]$ and $\mathsf{A}[\mathcal{V}_{pq}]$, respectively. Each augmented adjacency matrix has a corresponding augmented Laplacian matrix. The augmented Laplacian matrix of $S\subseteq\mathcal{V}$ is $\mathsf{L}[S]\triangleq\text{diag}(\mathsf{A}[S]\cdot 1_{N})-\mathsf{A}[S]\in\mathbb{R}^{N\times N}$. Consequently, the augmented Laplacian matrix of cluster $p$ and inter-cluster $pq$ are $\mathsf{L}[\mathcal{V}_{p}]$ and $\mathsf{L}[\mathcal{V}_{pq}]$, respectively.
For convenience, we introduce and henceforth use an alternative labeling scheme for the inter-clusters of $\mathcal{G}$ as determined by a cluster set $\mathcal{C}(\mathcal{V})$. Let $\mathtt{b}\colon[M]\times[M]\to[M^{\ast}]$ be a bijection. Given a cluster set $\mathcal{C}(\mathcal{V})$ with cardinality $M$, there are exactly $M^{\ast}$ pairs of distinct clusters with an inter-cluster. Hence, for some $p,q\in[M]$ with $p\neq q$, there exists a unique $r\in[M^{\ast}]$ such that $\mathtt{b}(p,q)=r$. By inductively using the bijection $\mathtt{b}$, let $\mathcal{V}^{r}\triangleq\mathcal{V}_{pq}$ for all $p,q\in[M]$ with $p\neq q$ such that $\mathtt{b}(p,q)=r$. Thus, the sub-graph and augmented Laplacian matrix of inter-cluster $pq$, for example, can also be written as $\mathcal{G}[\mathcal{V}^{r}]$ and $\mathsf{L}[\mathcal{V}^{r}]$, respectively. A static, undirected, and connected graph $\mathcal{G}$ with cluster and inter-cluster sub-graphs determined by cluster set $\mathcal{C}(\mathcal{V})$ and inter-cluster set $\mathcal{I}(\mathcal{V})$, respectively, is henceforth referred to as clustered network.
For each $p\in\mathcal{V}$, the neighbor set of agent $p$, that is, $\mathcal{N}_{p}$, can be partitioned using the cluster set $\mathcal{C}(\mathcal{V})$ and inter-cluster set $\mathcal{I}(\mathcal{V})$. The partition of $\mathcal{N}_{p}$ is given by $\mathsf{N}_{p}\triangleq\{\mathcal{N}_{p}\cap\mathcal{C}_{p},\ \mathcal{N}_{p}\cap\mathcal{V}^{1},\ \mathcal{N}_{p}\cap\mathcal{V}^{2},...,\mathcal{N}_{p}\cap\mathcal{V}^{M^{\ast}}\}$, where $S_{1}\cap S_{2}=\varnothing$ for any distinct $S_{1},S_{2}\in\mathsf{N}_{p}$ and
$\mathcal{N}_{p}=(\mathcal{N}_{p}\cap\mathcal{C}_{p})\bigcup\big{(}\cup_{r\in[M^{\ast}]}(\mathcal{N}_{p}\cap\mathcal{V}^{r})\big{)}.$ (3)
@@ II-D Hybrid Systems
:$\mathcal{H}\colon\Bigg{\{}\begin{aligned} \dot{\xi}&=f(\xi),\ &\xi\in C\phantom{,}\\ \xi^{+}&\in G(\xi),\ &\xi\in D,\end{aligned}$
where $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ denotes the flow map, $C\subset\mathbb{R}^{n}$ denotes the flow set, $G:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ denotes the jump map, and $D\subset\mathbb{R}^{n}$ denotes the jump set. A set $\Omega\subset\mathbb{R}_{\geq 0}\times\mathbb{Z}_{\geq 0}$ is a hybrid time domain, if there is a non-decreasing sequence of non-negative reals $(t_{j})_{j=0}^{m}$ with $m\in\mathbb{Z}_{\geq 0}\cup\{\infty\}$, $t_{0}=0$, and $t_{m}\in\mathbb{R}_{\geq 0}\cup\{\infty\}$, such that $\Omega=\cup_{j=1}^{m}(I_{j}\times\{j-1\})$, where all the intervals $I_{j}$, for $j<m$, are of the form $[t_{j-1},t_{j}]$, and $I_{m}$ is of the form $[t_{m-1},t_{m}]$ or $[t_{m-1},t_{m})$ when $m<\infty$. Note that $t_{m}=\infty$ is allowed when $m<\infty$; for $m=\infty$ there is no $t_{m}$. The time $t_{j}$ indicates the $j^{\text{th}}$ instant the state $\xi$ jumps, where the value of $\xi$ after a jump is denoted by $\xi^{+}$. A hybrid arc $\phi$ is a function $\phi:\text{dom }\phi\to\mathbb{R}^{n}$, such that dom $\text{dom }\phi\subset\mathbb{R}_{\geq 0}\times\mathbb{Z}_{\geq 0}$ is a hybrid time domain with jump sequence $(t_{j})_{j=0}^{m}$, and $\phi$ is a locally absolutely continuous function on $I_{j}$, for every $j$. A solution of $\mathcal{H}$ is a hybrid arc $\phi$ such that, for all $j>0$, $\phi(t,j-1)\in C$ and $\dot{\phi}(t,j-1)=f(\phi(t,j-1))$ for almost all $t\in I_{j}$; and $\phi(t_{j-1},j-1)\in D$ and $\phi(t_{j-1},j)\in G(\phi(t_{j-1},j-1))$. A solution $\phi$ to $\mathcal{H}$ is called maximal if $\phi$ cannot be extended, i.e., there does not exist a distinct solution $\psi$ to $\mathcal{H}$ such that dom $\text{dom }\phi\subseteq\text{dom }\psi$ and $\phi(t,j)=\psi(t,j)$ for all $(t,j)\in\text{dom }\phi$. A solution $\phi$ is called complete if dom $\text{dom }\phi$ is unbounded. A hybrid system $\mathcal{H}$ with data $(C,f,D,G)$ is said to satisfy the hybrid basic conditions if it satisfies [32, Assumption 6.5]. The following definition originates from [33, Section II].
@@
Let $\mathcal{H}$ be a hybrid system on $\mathcal{X}\subset\mathbb{R}^{n}$ with data $(C,f,D,G)$. A closed set $\mathcal{A}\subset\mathcal{X}$ is said to be globally exponentially stable (GES) for the hybrid system $\mathcal{H}$ if there exist constants $\kappa,\alpha\in\mathbb{R}_{>0}$ such that every maximal solution $\phi$ of $\mathcal{H}$ is complete and satisfies
@@
When considering perturbations on $\mathcal{H}$, input-to-state stability (ISS) is a useful analysis tool. The following definition comes from [34, Definition 2.2].
@@
The hybrid system with state $z$ and input $u$ given by
$\dot{z}$ $=f(z,u),$ $(z,u)\in C$ (4)
$z^{+}$ $\in G(z,u),$ $(z,u)\in D$

### v3

@@ 2.2 Graphs
Let $\mathcal{G}\coloneqq(\mathcal{V},\mathcal{E})$ be a graph on $N\in\mathbb{Z}_{\geq 2}$ nodes, where $\mathcal{V}\coloneqq[N]$ denotes the node set and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denotes the edge set. If $(p,q)\in\mathcal{E}$ implies $(q,p)\in\mathcal{E}$ for all distinct nodes $p,q\in\mathcal{V}$, then the graph $\mathcal{G}$ is said to be undirected. A path exists between nodes $p,q\in\mathcal{V}$ if there is a sequence of distinct nodes such that $(v_{0}=p,...,v_{k}=q)$ for $k\in\mathbb{Z}_{\geq 0}$, $(v_{s-1},v_{s})\in\mathcal{E}$, and $s\in[k]$. The graph $\mathcal{G}$ is said to be connected if there is a path joining any two distinct nodes in $\mathcal{V}$. The neighbor set of node $p$ is denoted by $\mathcal{N}_{p}\coloneqq\{q\in\mathcal{V}\setminus\{p\}:(p,q)\in\mathcal{E}\}$. The adjacency matrix of $\mathcal{G}$ is denoted by $\mathsf{A}\coloneqq[a_{pq}]\in\mathbb{R}^{N\times N}$, where $a_{pq}=1$ if and only if $(p,q)\in\mathcal{E}$, and $a_{pq}=0$ otherwise. Self-edges are not employed in this work, that is, $a_{pp}\coloneqq 0$ for all $p\in\mathcal{V}$. The degree matrix of $\mathcal{G}$ is denoted by $\Delta\coloneqq\text{diag}(\mathsf{A}\cdot 1_{N})\in\mathbb{R}^{N\times N}$. The Laplacian matrix of $\mathcal{G}$ is denoted by $\mathsf{L}\coloneqq\Delta-\mathsf{A}\in\mathbb{R}^{N\times N}$. The following result enables the stability analysis provided in Section 5.
@@
If $\mathcal{G}$ is static, undirected, and connected, then there exists an orthonormal basis $\beta\coloneqq\{v_{1},v_{2},...,v_{N}\}\subset\mathbb{R}^{N}$ for $\text{Range}(\mathsf{L})$ such that $v_{1}=(\sqrt{N}/N)1_{N}$. Consider the matrix $\mathsf{V}\coloneqq[v_{2},v_{3},...,v_{N}]\in\mathbb{R}^{N\times N-1}$ and projection $\mathsf{S}\coloneqq I_{N}-1_{N}1_{N}^{\top}/N\in\mathbb{R}^{N\times N}$. Then,
@@
The matrix $\mathsf{S}$ is a projection whose image is the orthogonal complement of the agreement subspace $\text{Span}(1_{N})$. That is, for any vector $\in\mathbb{R}^{N}, zz$ may be decomposed into the sum of two orthogonal vectors, i.e., $z=z^{\|}+z^{\perp}$, where $z^{\|}\coloneqq(1_{N}1_{N}^{\top}/N)z$ and $z^{\perp}\coloneqq\mathsf{S}z$. In addition, for any symmetric Laplacian matrix $\mathcal{L}$, one can show by direct computation that $\mathsf{S}\mathcal{L}=\mathcal{L}=\mathcal{L}\mathsf{S}$.
@@
2.3 Cluster and Inter-Cluster Sub-Graphs
@@
In this section, we introduce the precise notions of cluster and inter-cluster sub-graphs leading to the clustered networks discussed in this work. Please see [Pham.Messai.ea2020a, Pham.Messai.ea2020b, Shi.Hu.ea2023] for additional information on clustered networks. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a static, undirected, and connected graph. Let $\mathcal{C}(\mathcal{V})\coloneqq\{\mathcal{V}_{1},\mathcal{V}_{2},...,\mathcal{V}_{M}\}$ be a fixed partition of the set $\mathcal{V}$, where $M\in[N]$ is some constant. Consequently, $\mathcal{V}_{p}\subseteq\mathcal{V}$ for each $p\in[M]$, $\mathcal{V}_{p}\cap\mathcal{V}_{q}=\varnothing$ for each distinct pair $p,q\in[M]$, and $\cup_{p\in[M]}\mathcal{V}_{p}=\mathcal{V}$. The sets $\mathcal{C}(\mathcal{V})$ and $\mathcal{V}_{q}$, for any $q\in\mathcal{V}$, are referred to as the cluster set of $\mathcal{G}$ and cluster $q$, respectively. Since $\mathcal{C}(\mathcal{V})$ is an arbitrary partition of $\mathcal{V}$, for every $p\in[M]$, cluster $p$ contains at least one node, i.e., $|\mathcal{V}_{p}|\geq 1$. Each cluster induces a sub-graph of $\mathcal{G}$, i.e., for each $p\in[M]$, the induced sub-graph of cluster $p$ is given by $\mathcal{G}[\mathcal{V}_{p}]\coloneqq(\mathcal{V}_{p},\mathcal{E}[\mathcal{V}_{p}])$, such that $\mathcal{E}[\mathcal{V}_{p}]\coloneqq\{(r,s)\in\mathcal{E}\colon r,s\in\mathcal{V}_{p}\}$. Since $\mathcal{G}$ is undirected, $\mathcal{G}[\mathcal{V}_{p}]$ is undirected. In this work, we only consider cluster sets $\mathcal{C}(\mathcal{V})$ where, for each $p\in[M]$, $\mathcal{G}[\mathcal{V}_{p}]$ is connected. Moreover, since $\mathcal{C}(\mathcal{V})$ is a fixed partition of $\mathcal{V}$, for each $p\in\mathcal{V}$, node $p$ is assigned to one cluster for all time. In particular, for each $p\in\mathcal{V}$, let $\mathcal{C}_{p}$ denote the cluster to which node $p$ belongs, i.e., $p\in\mathcal{C}_{p}$ and $\mathcal{C}_{p}=\mathcal{V}_{q}$ for some $\mathcal{V}_{q}\in\mathcal{C}(\mathcal{V})$. The set of neighbors of node $p$ within $\mathcal{C}_{p}$ is, therefore, given by $\mathcal{N}_{p}\cap\mathcal{C}_{p}$.
@@
For any distinct pair $p,q\in[M]$, let $\mathcal{V}_{pq}\coloneqq\mathcal{V}(p,q)\cup\mathcal{V}(q,p)$, such that $\mathcal{V}(p,q)\coloneqq\{r\in\mathcal{V}_{p}\colon\exists_{s\in\mathcal{V}_{q}}\ (r,s)\in\mathcal{E}\}$. If $\mathcal{V}_{pq}\neq\varnothing$, then the set $\mathcal{V}_{pq}$ is referred to as inter-cluster $pq$. To simplify the development, we only consider the case where there is at most one inter-cluster between each pair of distinct clusters. The inter-cluster set of $\mathcal{G}$, as determined by $\mathcal{C}(\mathcal{V})$, is denoted by $\mathcal{I}(\mathcal{V})\coloneqq\{\mathcal{V}_{pq}\subset\mathcal{V}\colon\forall_{p\neq q\in\mathcal{V}}\ \mathcal{V}_{pq}\neq\varnothing\}$. Since $|\mathcal{C}(\mathcal{V})|=M$ and there can be at most one inter-cluster between each distinct pair of clusters, $|\mathcal{I}(\mathcal{V})|=M^{\ast}\in\{0,1,...,M(M-1)/2\}$. For each distinct pair of $p,q\in[M]$, the sub-graph of $\mathcal{G}$ induced by inter-cluster $pq$ is denoted by $\mathcal{G}[\mathcal{V}_{pq}]\coloneqq(\mathcal{V}_{pq},\mathcal{E}_{pq})$, where $\mathcal{E}_{pq}\coloneqq\mathcal{E}[\mathcal{V}_{pq}]\setminus(\mathcal{E}[\mathcal{V}_{p}]\cup\mathcal{E}[\mathcal{V}_{q}])$. Further, $\mathcal{G}[\mathcal{V}_{pq}]$ is undirected by construction. Figure 1 illustrates an example of a graph $\mathcal{G}$ that is divided into cluster and inter-cluster sub-graphs. Since $\mathcal{C}(\mathcal{V})$ and $\mathcal{G}$ are fixed, one has that $\mathcal{I}(\mathcal{V})$ is fixed as well. Thus, if node $p\in\mathcal{V}$ belongs to an inter-cluster, then it does so for all time. Moreover, it is possible for a node to belong to multiple inter-clusters simultaneously. The set of neighbors of node $p$ in inter-cluster $rs$ is, therefore, given by $\mathcal{N}_{p}\cap\mathcal{V}_{rs}$.
@@
To facilitate the analysis, $N\times N$ augmented adjacency and Laplacian matrices for a general sub-graph of $\mathcal{G}$ are introduced next. For any non-empty set $S\subseteq\mathcal{V}$, let $\mathsf{A}[S]\coloneqq[\mathsf{a}_{pq}]\in\mathbb{R}^{N\times N}$ with
$\mathsf{a}_{pq}\coloneqq\Bigg\{\begin{aligned} &1,&(p,q)\in\mathcal{E}[S]\\ &0,&\text{otherwise}\end{aligned}$
denote the augmented adjacency matrix of $\mathcal{G}[S]$. Therefore, the augmented adjacency matrix of cluster $p$ and inter-cluster $pq$ are $\mathsf{A}[\mathcal{V}_{p}]$ and $\mathsf{A}[\mathcal{V}_{pq}]$, respectively. Each augmented adjacency matrix has a corresponding augmented Laplacian matrix. The augmented Laplacian matrix of $S\subseteq\mathcal{V}$ is $\mathsf{L}[S]\coloneqq\text{diag}(\mathsf{A}[S]\cdot 1_{N})-\mathsf{A}[S]\in\mathbb{R}^{N\times N}$. Consequently, the augmented Laplacian matrix of cluster $p$ and inter-cluster $pq$ are $\mathsf{L}[\mathcal{V}_{p}]$ and $\mathsf{L}[\mathcal{V}_{pq}]$, respectively.
For convenience, we introduce and henceforth use an alternative labeling scheme for the inter-clusters of $\mathcal{G}$ as determined by a cluster set $\mathcal{C}(\mathcal{V})$. Let $\mathtt{b}\colon[M]\times[M]\to[M^{\ast}]$ be a bijection. Given a cluster set $\mathcal{C}(\mathcal{V})$ with cardinality $M$, there are exactly $M^{\ast}$ pairs of distinct clusters with an inter-cluster. Hence, for some $p,q\in[M]$ with $p\neq q$, there exists a unique $r\in[M^{\ast}]$ such that $\mathtt{b}(p,q)=r$. By inductively using the bijection $\mathtt{b}$, let $\mathcal{V}^{r}\coloneqq\mathcal{V}_{pq}$ for all $p,q\in[M]$ with $p\neq q$ such that $\mathtt{b}(p,q)=r$. Thus, the sub-graph and augmented Laplacian matrix of inter-cluster $pq$, for example, can also be written as $\mathcal{G}[\mathcal{V}^{r}]$ and $\mathsf{L}[\mathcal{V}^{r}]$, respectively. A static, undirected, and connected graph $\mathcal{G}$ with cluster and inter-cluster sub-graphs determined by cluster set $\mathcal{C}(\mathcal{V})$ and inter-cluster set $\mathcal{I}(\mathcal{V})$, respectively, is henceforth referred to as clustered network.
For each $p\in\mathcal{V}$, the neighbor set of agent $p$, that is, $\mathcal{N}_{p}$, can be partitioned using the cluster set $\mathcal{C}(\mathcal{V})$ and inter-cluster set $\mathcal{I}(\mathcal{V})$. The partition of $\mathcal{N}_{p}$ is given by $\mathsf{N}_{p}\coloneqq\{\mathcal{N}_{p}\cap\mathcal{C}_{p},\ \mathcal{N}_{p}\cap\mathcal{V}^{1},\ \mathcal{N}_{p}\cap\mathcal{V}^{2},...,\mathcal{N}_{p}\cap\mathcal{V}^{M^{\ast}}\}$, where $S_{1}\cap S_{2}=\varnothing$ for any distinct $S_{1},S_{2}\in\mathsf{N}_{p}$ and
$\mathcal{N}_{p}=(\mathcal{N}_{p}\cap\mathcal{C}_{p})\cup\big(\cup_{r\in[M^{\ast}]}(\mathcal{N}_{p}\cap\mathcal{V}^{r})\big).$ (3)
@@ 2.4 Hybrid Systems
$\mathcal{H}\colon\Bigg\{\begin{aligned} \dot{\xi}&=f(\xi),\ &\xi\in C\phantom{,}\\ \xi^{+}&\in G(\xi),\ &\xi\in D,\end{aligned}$
where $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ denotes the flow map, $C\subset\mathbb{R}^{n}$ denotes the flow set, $G:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ denotes the jump map, and $D\subset\mathbb{R}^{n}$ denotes the jump set. [^3] Within the context of multi-agent systems, the hybrid system $\mathcal{H}$ will be used to model the continuous-time and discrete-time behaviors of the ensemble. In particular, the flow set $C$ and jump set $D$ will be used to determine when the state of the ensemble, namely, $\xi$, may flow according to the flow map $f$ and/or jump according to the jump map $G$, respectively. For example, if the state of the ensemble is in the jump set, i.e., $\xi\in D$, then the state variable of the ensemble may jump according to $\xi^{+}\in G(\xi)$, where some variables in $\xi$ may change discretely while other variables in $\xi$ may be mapped to their current value under an identity transformation (since they may only change in continuous time). A set $\Omega\subset\mathbb{R}_{\geq 0}\times\mathbb{Z}_{\geq 0}$ is a hybrid time domain, if there is a non-decreasing sequence of non-negative reals $(t_{j})_{j=0}^{m}$ with $m\in\mathbb{Z}_{\geq 0}\cup\{\infty\}$, $t_{0}=0$, and $t_{m}\in\mathbb{R}_{\geq 0}\cup\{\infty\}$, such that $\Omega=\cup_{j=1}^{m}(I_{j}\times\{j-1\})$, where all the intervals $I_{j}$, for $j<m$, are of the form $t_{j-1},t_{j}]$, and $I_{m}$ is of the form $t_{m-1},t_{m}]$ or $t_{m-1},t_{m})$ when $m<\infty$. Note that $t_{m}=\infty$ is allowed when $m<\infty$; for $m=\infty$ there is no $t_{m}$. The time $t_{j}$ indicates the $j^{\text{th}}$ instant the state $\xi$ jumps, where the value of $\xi$ after a jump is denoted by $\xi^{+}$. A hybrid arc $\phi$ is a function $\phi:\text{dom }\phi\to\mathbb{R}^{n}$, such that $\text{dom }\phi\subset\mathbb{R}_{\geq 0}\times\mathbb{Z}_{\geq 0}$ is a hybrid time domain with jump sequence $(t_{j})_{j=0}^{m}$, and $\phi$ is a locally absolutely continuous function on $I_{j}$, for every $j$. A solution of $\mathcal{H}$ is a hybrid arc $\phi$ such that, for all $j>0$, $\phi(t,j-1)\in C$ and $\dot{\phi}(t,j-1)=f(\phi(t,j-1))$ for almost all $t\in I_{j}$; and $\phi(t_{j-1},j-1)\in D$ and $\phi(t_{j-1},j)\in G(\phi(t_{j-1},j-1))$. A solution $\phi$ to $\mathcal{H}$ is called maximal if $\phi$ cannot be extended, i.e., there does not exist a distinct solution $\psi$ to $\mathcal{H}$ such that $\text{dom }\phi\subseteq\text{dom }\psi$ and $\phi(t,j)=\psi(t,j)$ for all $(t,j)\in\text{dom }\phi$. A solution $\phi$ is called complete if $\text{dom }\phi$ is unbounded. A hybrid system $\mathcal{H}$ with data $(C,f,D,G)$ is said to satisfy the hybrid basic conditions if it satisfies [Goebel.Sanfelice.ea2012, Assumption 6.5]. The following definition originates from [Teel.Forni.ea2013, Section II].
@@
Let $\mathcal{H}$ be a hybrid system on $\mathcal{X}\subset\mathbb{R}^{n}$ with data $(C,f,D,G)$. A closed set $\mathcal{A}\subset\mathcal{X}$ is said to be globally exponentially stable for the hybrid system $\mathcal{H}$ if there exist constants $\kappa,\alpha\in\mathbb{R}_{>0}$ such that every maximal solution $\phi$ of $\mathcal{H}$ is complete and satisfies
@@
When considering perturbations on $\mathcal{H}$, input-to-state stability offers a convenient tool for analysis. The following definition is a slight modification of [Li.Phillips.ea2018, Definition 2.2].
@@
The hybrid system with state $z$ and perturbation d{d} given by
$\dot{z}$ $=f(z,{d}),\$$(z,{d})\in C$ (4)
$z^{+}$ $\in G(z,{d}),\$$(z,{d})\in D$